Hello .. suppose i have inflation rate f = 3% compounded annually and i want to use it in monthly payment .. how i can do that?


please show all steps
Answer #1

In order to use the inflation rate for monthly payments we will have to find the monthly equivalent rate of the annually compounded inflation rate.

Let the monthly compounded rate be m%

The following formula will be used.

(1 + annually compounded rate) = (1 + monthly compounded rate)12

or, (1+ f) = (1 + m)12

or, m = {(1+ f)1/12 - 1 }

or, m = (1.031/12 - 1)

or, m = 0.2466%

Kindly up vote my answer.